OpenStudy (anonymous):

Solve: 8t - t + 1 = 10 - 2t

OpenStudy (anonymous):

\(\ \Huge 8t-t+1=10-2t \rightarrow 8t-t=9-2t \). \(\ \Huge 8t-t=9-2t \rightarrow 7t=9-2t \rightarrow 9t=9, \) \(\ \Huge So, t=1. \)